If y is directly proportional to x and inversely proportional to the square of z, and if y=0.384 when x=6 and z=5, find the constant of variation.
Answer by stanbon(75887)

If y is directly proportional to x and inversely proportional to the square of z, and if y=0.384 when x=6 and z=5, find the constant of variation.
--------
y = kx/z^2
--------
0.384 = k*6/(5^2)
k = 0.384*25/6
k = 1.6
===========
Cheers,
Stan H.